A spring-mass system has a constant force of 180 N/m. What mass will make it to vibrate at : a) 3.0 Hz and b) 5 Hz?


1
Expert's answer
2021-01-26T08:17:05-0500

a)


"f=\frac{1}{2\pi}\sqrt{\frac{k}{m}}\\3=\frac{1}{2\pi}\sqrt{\frac{180}{m}}\\m=0.51\ kg"

b)


"f=\frac{1}{2\pi}\sqrt{\frac{k}{m}}\\5=\frac{1}{2\pi}\sqrt{\frac{180}{m}}\\m=0.18\ kg"